OpenStudy (anonymous):

What is the solution to the equation? -4x 5(x-2)=-16 10 x=??

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

Is this -4x+5(x-2)=-16+10? Some of the signs are missing from the equation you put down.

OpenStudy (anonymous):

yes. for some reason my + marks dont work in a question

OpenStudy (anonymous):

That's okay. -4x+5(x-2)=-16+10 The first step is to distribute. -4x + 5x - 10 = -16 + 10 Combine like terms. x-10=-6 Add 10 to both sides. x=4 You can check the solution by plugging it into the original equation.

OpenStudy (anonymous):

Thank you! I have some more. Can you help me @verlieren

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

x-7.6+9.4=-8.4

OpenStudy (anonymous):

@verlieren

OpenStudy (anonymous):

The first thing you need to do is combine like terms. What are the like terms that you see on the left side?

OpenStudy (anonymous):

7.6 and 9.4

OpenStudy (anonymous):

Don't forget the negative sign in front of 7.6. It would really be -7.6 + 9.4. -7.6+9.4=1.8, so you now have x+1.8=-8.4. The goal of every equation is to isolate the variable, so what should you do with 1.8 now?

OpenStudy (anonymous):

That's a minus sign im pretty sure

OpenStudy (anonymous):

@verlieren

OpenStudy (anonymous):

A minus sign is the same thing as saying x+(-7.6). Understanding this is huge because this kind of thing comes up a lot in algebra. To demonstrate this, let's pretend it really is 7.6 + 9.4 rather than -7.6 + 9.4: x-7.6+9.4=-8.4 x-17=-8.4 x=8.6 Let's plug that "solution" into the original equation. 8.6-7.6+9.4=-8.4 10.4=-8.4 That isn't true, right? So it does have to be -7.6 + 9.4. x+1.8=-8.4 Subtract 1.8 from each side to isolate x. x=-10.2 Let's check our work. -10.2-7.6+9.4=-8.4 -8.4=-8.4 There you have it! Make sure you get this down. It's crucial to your success in Algebra 1 (if you're not already taking it) and Algebra 2.